OpenStudy (anonymous):

When the cost of a car wash was $8, there were 90 customers. When the cost was reduced to $4, there were 170 customers. Assume the relationship is linear.

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

What's the question? It just says to assume the relationship is linear..

OpenStudy (igreen):

You have two points on a graph. (8, 90) and (4, 170).

OpenStudy (anonymous):

A line models the relationship between the cost, x, and the number of customers, y. What is the slope of the line?

OpenStudy (anonymous):

^^^ the Question

OpenStudy (igreen):

You have to find the slope between (8, 90) and (4, 170). To do that you have to use the slope formula, \(m = \dfrac{y_2-y_1}{x_2-x_1}\). In this case: \(y_2 = 170\) \(y_1 = 90\) \(x_2 = 4\) \(x_1 = 8\) Plug them in: \(m = \dfrac{170-90}{4-8}\) Subtract: \(m = \dfrac{80}{-4}\) \(m = -20\) So the slope is -20.